# Condensed-space methods for nonlinear programming on GPUs

### v1

@@ Abstract
This paper explores two condensed-space interior-point methods to efficiently solve large-scale nonlinear programs on graphics processing units (GPUs). The interior-point method solves a sequence of symmetric indefinite linear systems, or Karush-Kuhn-Tucker (KKT) systems, which become increasingly ill-conditioned as we approach the solution. Solving a KKT system with traditional sparse factorization methods involve numerical pivoting, making parallelization difficult. A solution is to condense the KKT system into a symmetric positive-definite matrix and solve it with a Cholesky factorization, stable without pivoting. Although condensed KKT systems are more prone to ill-conditioning than the original ones, they exhibit structured ill-conditioning that mitigates the loss of accuracy. This paper compares the benefits of two recent condensed-space interior-point methods, HyKKT and LiftedKKT. We implement the two methods on GPUs using MadNLP.jl, an optimization solver interfaced with the NVIDIA sparse linear solver cuDSS and with the GPU-accelerated modeler ExaModels.jl. Our experiments on the PGLIB and the COPS benchmarks reveal that GPUs can attain up to a tenfold speed increase compared to CPUs when solving large-scale instances.
@@ 1 Introduction
Graphics processing units (GPUs) are driving the advancement of scientific computing, their most remarkable success being the capabilities to train and utilize large artificial intelligence (AI) models. GPUs offer two practical advantages: (1) massive parallel computing capability for applications that can exploit coarse-grain parallelism and high-memory bandwidth and (2) power efficiency due to requiring fewer transistors to process multiple tasks in parallel utilizing “single instruction, multiple data” (SIMD) parallelism.
While GPUs have made significant strides in enhancing machine learning applications, their adoption in the mathematical programming community has been relatively limited. This limitation stems primarily from the fact that most optimization solvers were developed in the 1990s and are heavily optimized for CPU architectures. Additionally, the utilization of GPUs has been impeded by the challenges associated with sparse matrix factorization routines, which are inherently difficult to parallelize on SIMD architectures. Nevertheless, recent years have witnessed notable advancements that are reshaping this landscape.
@@
    Improved sparse matrix operations: The performance of sparse matrix operations has seen substantial improvements in the CUDA library, largely attributed to the integration of novel tensor cores in recent GPUs markidis2018nvidia .
@@
    Interest in batch optimization: There is a growing interest in solving parametric optimization problems in batch mode, for problems sharing the same structure but with different parameters amos2017optnet ; pineda2022theseus .
@@
    Advancements in automatic differentiation: GPUs offer unparalleled performance for automatic differentiation, benefiting both machine learning jax2018github and scientific computing applications enzyme2021 . Engineering problems often exhibit recurring patterns throughout the model. Once these patterns are identified, they can be evaluated in parallel within a SIMD framework, enabling near speed-of-light performance shin2023accelerating .
@@
    Role in exascale computing: With the emergence of new exascale supercomputers (e.g., Frontier and Aurora), the capabilities to run on GPUs have become central for supercomputing.
1.1 Solving optimization problems on GPU: current state-of-the-art
For all the reasons listed before, there is an increasing interest to solve optimization problems on GPUs. We now summarize the previous work on solving classical—large-scale, sparse, constrained—mathematical programs on GPUs.
GPU for mathematical programming.
The factorization of sparse matrices encountered within second-order optimization algorithms has been considered to be challenging on GPUs. For this reason, practitioners often has resorted to using first-order methods on GPUs, leveraging level-1 and level-2 BLAS operations that are more amenable to parallel computations. First-order algorithms depend mostly on (sparse) matrix-vector operations that run very efficiently on modern GPUs. Hence, we can counterbalance the relative inaccuracy of the first-order method by running more iterations of the algorithm. A recent breakthrough lu2023cupdlp ; lu2023cupdlp2 based on the primal-dual hybrid gradient method has demonstrated that a first-order algorithm can surpass the performance of Gurobi, a commercial solver, in tackling large-scale linear programs. This performance gain is made possible by executing the first-order iterations solely on the GPU through an optimized codebase.
GPU for batched optimization solvers.
The machine learning community has been a strong advocate for porting mathematical optimization on the GPU. One of the most promising applications is embedding mathematical programs inside neural networks, a task that requires batching the solution of the optimization model for the training algorithm to be efficient amos2017optnet ; pineda2022theseus . This has led to the development of prototype code solving thousands of (small) optimization problems in parallel on the GPU. Furthermore, batched optimization solvers can be leveraged in decomposition algorithms, when the subproblems share the same structure kimLeveragingGPUBatching2021 . However, it is not trivial to adapt such code to solve large-scale optimization problems, as the previous prototypes are reliant on dense linear solvers to compute the descent directions.
GPU for nonlinear programming.
The success of first-order algorithms in classical mathematical programming relies on the convexity of the problem. Thus, this approach is nontrivial to replicate in nonlinear programming: Most engineering problems encode complex physical equations that are likely to break any convex structure in the problem. Previous experiments on the alternating current (AC) optimal power flow (OPF) problem have shown that even a simple algorithm as the alternating direction method of multipliers (ADMM) has trouble converging as soon as the convergence tolerance is set below $10^{-3}$ kimLeveragingGPUBatching2021 .
Thus, second-order methods remain a competitive option, particularly for scenarios that demand higher levels of accuracy and robust convergence. Second-order algorithms solve a Newton step at each iteration, an operation relying on non-trivial sparse linear algebra operations. The previous generation of GPU-accelerated sparse linear solvers were lagging behind their CPU equivalents, as illustrated in subsequent surveys swirydowicz2021linear ; tasseff2019exploring . Fortunately, sparse solvers on GPUs are becoming increasingly better: NVIDIA has released in November 2023 the cuDSS sparse direct solver that implements different sparse factorization routines with remarkably improved performance. Our benchmark results indicate that cuDSS is significantly faster than the previous sparse solvers using NVIDIA GPUs (e.g., published in shin2023accelerating ). Furthermore, variants of interior point methods have been proposed that does not depend on numerical pivoting in the linear solves, opening the door to parallelized sparse solvers. Coupled with a GPU-accelerated automatic differentiation library and a sparse Cholesky solver, these nonlinear programming solvers can solve optimal power flow (OPF) problems 10x faster than state-of-the-art methods shin2023accelerating .
There exist a few alternatives to sparse linear solvers for solving the KKT systems on the GPU. On the one hand, iterative and Krylov methods depend only on matrix-vector products to solve linear systems. They often require non-trivial reformulation or specialized preconditioning of the KKT systems to mitigate the inherent ill-conditioning of the KKT matrices, which has limited their use within the interior-point methods curtisNoteImplementationInteriorpoint2012 ; rodriguezScalablePreconditioningBlockstructured2020 . New results are giving promising outlooks for convex problems ghannad2022linear , but nonconvex problems often need an Augmented Lagrangian reformulation to be tractable cao2016augmented ; regev2023hykkt . In particular, regev2023hykkt presents an interesting use of the Golub and Greif hybrid method golub2003solving to solve the KKT systems arising in the interior-point methods, with promising results on the GPU. On the other hand, null-space methods (also known as reduced Hessian methods) reduce the KKT system down to a dense matrix, a setting also favorable for GPUs. Our previous research has shown that the method plays nicely with the interior-point methods if the number of degrees of freedom in the problem remains relatively small pacaud2022condensed .
@@ 1.2 Contributions
In this article, we assess the current capabilities of modern GPUs to solve large-scale nonconvex nonlinear programs to optimality. We focus on the two condensed-space methods introduced respectively in regev2023hykkt ; shin2023accelerating . We re-use classical results from wright1998ill to show that for both methods, the condensed matrix exhibits structured ill-conditioning that limits the loss of accuracy in the descent direction (provided the interior-point algorithm satisfies some standard assumptions). We implement both algorithms inside the GPU-accelerated solver MadNLP, and leverage the GPU-accelerated automatic differentiation backend ExaModels shin2023accelerating . The interior-point algorithm runs entirely on the GPU, from the evaluation of the model (using ExaModels) to the solution of the KKT system (using a condensed-space method running on the GPU). We use CUDSS.jl Montoison_CUDSS , a Julia interface to the NVIDIA library cuDSS, to solve the condensed KKT systems. We evaluate the strengths and weaknesses of both methods, in terms of accuracy and runtime. Extending beyond the classical OPF instances examined in our previous work, we incorporate large-scale problems sourced from the COPS nonlinear benchmark dolan2004benchmarking . Our assessment involves comparing the performance achieved on the GPU with that of a state-of-the-art method executed on the CPU. The findings reveal that the condensed-space IPM enables a remarkable ten-time acceleration in solving large-scale OPF instances when utilizing the GPU. However, performance outcomes on the COPS benchmark exhibit more variability.
@@ 1.3 Notations
[…]
@@
  • We write $x=\Theta(h^{p})$ if there exists two constants $0<a<b$ such that $ah^{p}\leq\|x\|\leq bh^{p}$.

### v2

# Condensed Interior-Point Methods for Scalable Nonlinear Programming on GPUs

@@ Abstract
This paper explores two variants of condensed-space interior-point methods designed for GPUs—HyKKT and LiftedKKT—by analyzing their numerical properties through error analysis and assessing their real-world performance via extensive numerical experiments with a fully GPU-resident software implementation. Traditional implementations of interior-point methods (IPMs) involve solving indefinite augmented KKT systems repeatedly by utilizing direct sparse solvers based on the LBL factorization with sophisticated numerical pivoting strategies. While this method achieves high performance and robustness on CPUs, the serial nature of numerical pivoting presents challenges for effective implementation on GPUs. Recently, multiple condensed-space IPM strategies have emerged to address this issue by transforming the KKT system into a symmetric positive-definite matrix, which is more suitable for factorization on GPUs. However, these methods have been demonstrated only on optimal power flow instances, and their numerical properties are currently inadequately understood. In this study, we demonstrate that although the condensed systems show increased ill-conditioning, the inherent structures of the condensed KKT system effectively counterbalance potential accuracy loss in the IPM. Furthermore, we provide numerical results that thoroughly assess the capabilities of a fully GPU-resident nonlinear programming software stack, comprising MadNLP.jl (a filter line-search IPM solver), cuDSS (a direct sparse solver leveraging Cholesky factorization), and ExaModels.jl (a modeling framework), by benchmarking their performance against the pglib-opf and CUTEst libraries. Our findings suggest that the GPU framework holds promise for solving highly sparse large-scale nonlinear programs, such as optimal power flow instances, except for diminished robustness and limited speedups for edge cases observed within CUTEst instances.
@@ 1 Introduction
General-purpose GPU computing has emerged as a fundamental component of scientific computing and machine learning, particularly in its ability to train and deploy large-scale artificial intelligence (AI) models. As a computing platform, the modern GPU hardware offer two major advantages: (1) substantial parallel computing power and (2) improved power efficiency.
Despite their achievements in scientific computing and machine learning, GPUs have been underutilized in classical mathematical programming, including linear programming (LP) and nonlinear programming (NLP). In machine learning applications, the parallelization is mostly utilized at the low-level matrix operations, via the libraries like NVIDIA’s cuBLAS or AMD’s rocBLAS. In contrast, most classical mathematical programming applications exhibit limited parallelism at the BLAS operation level due to the inherent sparsity of the underlying problem instances, which prevents straightforward utilization of parallel resources. Although certain categories of sparse matrix operations can be parallelized, the development of efficient sparse matrix kernels on GPUs poses significant challenges, particularly for sparse matrix factorizations. Classical solution methodologies typically depend on direct linear solvers that employ sophisticated numerical pivoting strategies to maintain numerical stability. However, these techniques are inherently sequential, creating difficulties for effective parallelization on GPUs. As a result, until recently, the full potential of GPU computing has not been realized within the mathematical programming domain.
Recent advancements in GPU computing techniques are beginning to change this landscape. Here are some significant developments, driven either by the hardware, software, or the emergence of new applications:
@@
    Development of mature direct sparse linear solvers: Recent developments in GPU-based sparse direct solvers, such as NVIDIA’s cuDSS nvidiaNVIDIACuDSSPreview , have significantly improved the performance of the direct, repetitive solution of sparse linear systems. These solvers now offer performance that rivals or exceeds that of traditional CPU-based solvers, though the types of factorizations currently supported are limited to Cholesky, $\mathrm{LDL^{T}}$, and LU as of now nvidiaNVIDIACuDSSPreview ; shin2023accelerating .
@@
    Progress in automatic differentiation: GPUs are well-suited for automatic differentiation, a key component of both scientific simulations and machine learning jax2018github ; enzyme2021 . Efficient implementations for algebraic modeling and derivative evaluation on GPUs have now become available. For large-scale problem instances with highly repetitive patterns, the derivative evaluation can be performed at speeds that are orders of magnitude faster than their CPU counterparts shin2023accelerating . Notably, derivative evaluation kernels can be created in a fully automated fashion. This removes the barrier of having to write custom derivative evaluation kernels for each problem instance, which has been a major bottleneck in the past.
@@
    Rise of GPU-dominated HPC: With the emergence of exascale supercomputers (e.g., Frontier, Aurora), GPU support has become a core requirement for utilizing these HPC resources in mathematical optimization applications.
@@
    Growing interest in batched optimization: In machine learning, embedding optimization problems into neural networks requires solving many small problems in parallel. This has led to the development of GPU-native batched solvers capable of handling thousands of small problems simultaneously amos2017optnet ; pineda2022theseus . Similar ideas have been applied to decomposition methods in large-scale optimization kimLeveragingGPUBatching2021 , but extending these methods to large-scale problems remains difficult, as current implementations often rely on dense linear algebra.
Due to this renewed interest in GPU-based optimization, there has been a surge of research in this area, particularly in the context of developing scalable algorithms and software implementations that can harness the power of GPUs. Here, we provide a general overview of the state of the art in GPU-based optimization, highlighting the challenges and opportunities in this area. The recent developments in GPU-based optimization can be broadly categorized into three main approaches: (i) first-order methods, (ii) second-order methods with iterative linear algebra, and (iii) second-order methods with sparse direct solvers. In terms of numerical implementation, the first two approaches rely mostly on repetitive sparse matrix-vector multiplications, while the third approach requires sparse direct solvers.
First-order methods for large-scale optimization.
Although first-order methods have been extensively used for large-scale optimization, especially in machine learning applications, their application to classical mathematical programming remains limited due to their slow linear convergence rate. However, for extremely large-scale instances of linear programs, where the numerical factorization of the KKT systems or normal equations is infeasible, first-order methods can be a promising alternative. Recently, a special type of primal-dual first-order method, known as the primal-dual hybrid gradient (PDHG) method, has been developed to target such extremely large problems. This method is particularly effective for large-scale linear programs, as it can handle problems with billions of variables and constraints. This algorithm is implemented within a solver called PDLP, and it is adapted to GPU architectures lu2023cupdlp ; lu2023cupdlp2 . Recently, this algorithm has been enhanced with the restarted Halpern PDHG method with reflection and a PID-controlled primal weight update strategy, and has been shown to achieve greater speed luCuPDLPFurtherEnhanced2025 . Although these methods have faced challenges in obtaining high precision solutions, for extremely large-scale instances of linear programs, they have shown to outperform even commercial solvers like Gurobi when fully GPU-accelerated lu2023cupdlp ; lu2023cupdlp2 .
Second-order methods with iterative linear algebra.
An alternative to first-order methods is to perform second-order methods with iterative linear algebra, such as Krylov subspace methods. Krylov subspace methods require only matrix-vector products and are inherently parallelizable. These methods are well-suited for large-scale optimization problems, as they can efficiently solve the KKT systems arising in interior-point methods without requiring full matrix factorizations. However, most second-order algorithms, such as the primal-dual interior-point method, rely on barrier functions to handle inequality constraints, which can lead to ill-conditioned KKT systems; thus, the iterative linear solvers need to be employed with careful KKT system reformulation and/or preconditioning strategies curtisNoteImplementationInteriorpoint2012 ; rodriguezScalablePreconditioningBlockstructured2020 . Promising results have been reported in convex settings schubigerGPUAccelerationADMM2020 by combining the operator splitting methods and preconditioned conjugate gradient. For nonlinear programming, a strategy based on the Augmented Lagrangian method, interior-point method, and preconditioned conjugate gradient method has been proposed cao2016augmented .
Second-order methods for large-scale optimization.
The previous two strategies are well suited to large-scale convex optimization, but extending their success to general nonlinear programming remains challenging. For nonconvex settings, first-order or iterative linear algebra-based implementations often struggle to achieve high accuracy. For instance, it has been previously reported that ADMM fails to converge reliably below a tolerance of $10^{-3}$ on the AC optimal power flow (OPF) problem kimLeveragingGPUBatching2021 . However, second-order methods with direct linear algebra, when implemented correctly, can offer significantly faster convergence and more robust convergence behaviors. By solving a Newton system at each iteration, they can provide fast local convergence and high robustness, even in the presence of nonconvexities.
Second-order methods with direct linear algebra are substantially more challenging to implement on GPUs. This is because direct solution methods require the efficient factorization of large sparse, indefinite Karush–Kuhn–Tucker (KKT) systems. Even as of now, to the best of our knowledge, no GPU-based direct solver can directly handle these sparse indefinite KKT systems encountered in large-scale nonlinear optimization applications. Historically, GPU-based sparse solvers have lagged behind their CPU counterparts in both performance and robustness. Notably, it has been reported that GPU solvers do not offer significant acceleration for power system optimization problems swirydowicz2021linear ; tasseff2019exploring . Although several direct sparse solvers, such as CUSOLVERRF, were available on GPUs, they were not widely adopted in practice for solving large-scale nonlinear optimization problems due to their limited performance and robustness.
This situation has changed in recent years, with two advances: the development of pivoting-free interior-point methods and the emergence of new GPU-native sparse direct solvers. First, the KKT system reformulation strategy, known as condensed-space methods, has been proposed to address the challenges of solving indefinite KKT systems within interior-point methods. There are various forms of such reformulations, including the equality relaxation method (so-called LiftedKKT) and the hybrid Golub-Greif method (so-called HyKKT; adapted in regev2023hykkt ). Alternatively, null-space methods (also known as reduced-Hessian methods) eliminate equality constraints to reduce the KKT system to a dense problem, which maps well to GPU architectures pacaud2022condensed . These strategies transform the indefinite KKT system into a symmetric positive-definite (SPD) system, which can be solved using Cholesky factorization with static pivoting.
Furthermore, recent advances in GPU-based sparse direct solvers have significantly improved their performance and robustness. Especially for solving positive definite or quasi-definite systems, which do not require sophisticated pivoting strategies, GPU solvers now often surpass the performance of CPU-based solvers. Notably, in November 2023, NVIDIA released cuDSS, a new sparse direct solver providing efficient sparse factorizations with significantly improved performance over earlier GPU-based solvers. These advances, combined with recent developments in automatic differentiation and interior-point methods that avoid numerical pivoting, enable robust and scalable second-order methods to be implemented fully on GPUs. Coupling cuDSS with a GPU-native AD engine allows us to solve large-scale nonlinear problems like OPF up to 10× faster than previous state-of-the-art methods shin2023accelerating .
Nevertheless, the condensed-space methods have several known limitations. Although condensed KKT systems remain sparse in many problem instances, performing the elimination of the inequality constraints often leads to significantly increased nonzero entries in the condensed KKT system for certain instances. The worst case is a fully dense constraint where both the equality relaxation and HyKKT methods mentioned above can lead to fully dense condensed KKT systems, leading to a prohibitively expensive to solve. Due to this aspect, condensed-space methods are inherently unsuitable for problems with dense constraints. Accordingly, when these methods are compared against standard benchmark sets, such as CUTEst, they are bound to exhibit varied performance, due to the presence of the instances with such dense constraints. Furthermore, due to the nature of the condensed KKT systems, the condition number of the condensed KKT system can be significantly larger than that of the original KKT system, which can lead to numerical issues.
A full Software Stack for GPU-Accelerated Nonlinear Programming.
The state-of-the-art nonlinear optimization software stack consists of three key components: (i) a high-level modeling interface, (ii) a nonlinear optimization solver, (iii) a sparse linear solver. In the classical CPU-based setting, the most widely used software stack for nonlinear optimization consists of AMPL fourer1990ampl for modeling, Ipopt for the solver, and the HSL library (MA27, MA57, etc.) for sparse factorization. It is important to note that in order to fully benefit from the computational power of GPUs, the entire optimization pipeline must be implemented on the GPU. If any component is not GPU-compatible, it is necessary to introduce costly CPU–GPU communication, similar to double-buffering in MPI-based distributed systems. This avoids expensive transfers between host (CPU) and device (GPU) memory. For older generations of the PCIe interface, this can be a major bottleneck. Thus, in order to achieve high performance, it is essential that all components of the optimization pipeline are GPU-native and operate directly in GPU memory. The GPU-native software stack used within this paper is provided by MadNLP.jl, ExaModels.jl, and cuDSS. This paper focuses on the second component, the nonlinear optimization solver, while the other two components are briefly described below.
ExaModels.jl is an algebraic modeling system designed for efficient model function and derivative evaluations on GPUs shin2023accelerating . ExaModels.jl provides a symbolic modeling front-end where users can define their optimization problems in a structure-revealing fashion, similarly to the constraint template implemented in Gravity hijazi2018gravity . Unlike classical modeling interfaces such as AMPL, JuMP, and CasADi, which provide flexible modeling features, ExaModels.jl follows a different philosophy: it enforces strict compliance with the iterator-based modeling syntax, allowing the modeling system to store the model equations in a structured fashion. This ensures that the model function and derivative evaluations can be implemented as parallelizable GPU kernels. For the structured model functions, ExaModels.jl applies the reverse-mode automatic differentiation to compute the derivatives, then executed over GPU-resident data arrays. To the best of our knowledge, ExaModels.jl is currently the only available GPU-native sparse modeling interface that provides automatic differentiation capabilities for nonlinear programming beyond unconstrained problems in machine learning.
NVIDIA’s cuDSS library nvidiaNVIDIACuDSSPreview provides a high-performance sparse direct solver. The library supports Cholesky, LDL, and LU factorizations, enabling the efficient solution of symmetric positive definite and quasi-definite systems. When the hybrid host/device execution mode is disabled, reordering (a major part of the analysis phase) is executed on the host, while symbolic factorization (another part of the analysis phase), numerical factorization, and solve are executed on the GPU. Since ordering is inherently sequential, it is precomputed once on the CPU and transferred to the GPU without any impact on performance. Partial pivoting strategies are available, but the 2x2 pivoting strategy — commonly employed within sparse indefinite solvers – is not supported, thus limiting their direct usage within the classical augmented system-based interior-point methods. CUDSS.jl, a community-maintained Julia wrapper for cuDSS, provides a user-friendly interface to the library, enabling seamless integration with Julia-based optimization solvers such as MadNLP.jl.
@@ 1.1 Contributions
In this article, we aim to theoretically and empirically assess the capabilities of the GPU-accelerated solution strategies for large-scale sparse nonlinear programs. Although we utilize the full GPU-resident software stack, we focus on the nonlinear optimization solver component, which is based on filter line-search interior point method with two variants of condensed-space KKT system reformulation strategies: HyKKT regev2023hykkt and LiftedKKT shin2023accelerating . Our main contributions are summarized below:
  1. 1.
    We perform a rigorous error analysis of the condensed-space IPM methods for two condensed-space methods under study: HyKKT and LiftedKKT. We extend classical results from wright1998ill to show that the condensed KKT matrices are structurally ill-conditioned in a controlled manner. This yields accurate error bounds for the Schur complement used in the HyKKT method as well as for the condensed KKT system within LiftedKKT.
  2. 2.
    We provide a fully GPU-resident software stack—consisting of ExaModels.jl (derivative evaluation), MadNLP.jl (interior point solver), and NVIDIA cuDSS —for solving large-scale sparse nonlinear optimization problems, based on two different condensed-space KKT system reformulation strategies: HyKKT and LiftedKKT. Compared to our previous work shin2023accelerating , we (i) have additionally implemented the HyKKT method, and (ii) have replaced the linear solver cusolverRF used in shin2023accelerating with NVIDIA’s new cuDSS library. As a result, the interior-point method is executed entirely on the GPU, from model evaluation to KKT system solving.
  3. 3.
    We conduct a comparative study of the two condensed-space methods, analyzing their accuracy and runtime performance. We go beyond traditional OPF test cases by including large-scale problems from the CUTEst benchmark gould2015cutest . Our findings demonstrate that the condensed-space IPM enables a remarkable tenfold acceleration in solving large-scale OPF instances when utilizing the GPU. However, performance outcomes on the CUTEst benchmark exhibit more variability. For larger problems (where Ipopt takes longer than 1 second), our GPU solver is competitive with the CPU-based Ipopt solver, but exhibits less robustness and performance variability across different problem types. One of the main reasons for variability in performance is due to the fact that a single dense row can incur a dense condensed KKT system, which can make the factorization of the condensed KKT system significantly more expensive.
  4. 4.
    We demonstrate that condensed-space methods can be beneficial for a wide range of parallel architectures, not limited to GPUs. In fact, the condensed-space strategies presented in this paper are suitable not only for GPUs but also for other parallel architectures, such as multi-core CPUs on shared memory. We provide numerical results of the two condensed-space methods on multi-core CPUs, showing that they can offer substantial speed-ups for large-scale instances, though less pronounced than on GPUs. In particular, our numerical results reveal that the parallel CPU solvers, such as HSL MA86 and Pardiso, are particularly efficient at solving the condensed KKT systems. On the largest OPF instance, we observe that the parallel CPU solvers are almost 5 times faster than HSL MA27, whereas NVIDIA cuDSS is itself 25 times faster than HSL MA27.
@@ 1.2 Notations
By default, the norm $\|\cdot\|$ refers to the 2-norm. We define the conditioning of a matrix $A$ as $\kappa_{2}(A)=\|A\|\|A^{-1}\|$. For any real number $x$, we denote by $\widehat{x}$ its floating point representation. We denote $\mathbf{u}$ as the smallest positive number such that $\widehat{x}\leq(1+\tau)x$ for $|\tau|<\mathbf{u}$. In double precision, $\mathbf{u}=1.1\times 10^{-16}$. We use the following notations to proceed with our error analysis. For $p\in\mathbb{N}$ and a positive variable $h$:
@@
  • We write $x=\Theta(h^{p})$ if there exist two constants $0<a<b$ such that $ah^{p}\leq\|x\|\leq bh^{p}$.